OpenStudy (anonymous):

fine the difference quotient f(3+h)-f(3) and simplify if f(x) is x^2-2x+9

OpenStudy (anonymous):

divided by h

OpenStudy (anonymous):

did you find \(f(3)\) ? you need that first

OpenStudy (anonymous):

no im having dificulties

OpenStudy (anonymous):

it is a direct computation if you cannot find it, the other part will make less than no sense to you

OpenStudy (anonymous):

do you just plug 3 into x^2-2x+9?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

\[f(x) = x^2-2x+9\\ f(\clubsuit)=\clubsuit^2-2\clubsuit +9\\ f(3)=3^2-2\times 3+9\]\]

OpenStudy (anonymous):

let me know when you get \(12\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next you need to do something harder find \(f(3+h)\)

OpenStudy (anonymous):

ok i got 12

OpenStudy (anonymous):

\[f(\clubsuit)=\clubsuit^2-2\clubsuit +9\\\] where you see a \(\clubsuit\) replace it by \(3+h\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

im not sure if its right i got h^2+4h+12

OpenStudy (anonymous):

think so, let me check

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok now subtract \[f(3+h)-f(3)=h^2+4x+12-12=?\]

OpenStudy (xapproachesinfinity):

Good^_^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got h^2+4h

OpenStudy (anonymous):

ok good last step divide both terms by \(h\)

OpenStudy (anonymous):

h+4 :D?

OpenStudy (anonymous):

do you always find f(whatever number) and then the rest

OpenStudy (anonymous):

yes you are done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw hope it was more or less clear